OpenStudy (anonymous):

Under certain circumstances, the number of generations of "good" bacteria needed to increase the frequency of a particular gene from is n=4.572 integral 1/((x^2)(1-x)) dx

OpenStudy (anonymous):

\[n=4.572\int\limits_{0.1}^{0.6}1/(x^2(1-x))dx\]

OpenStudy (anonymous):

find n, rounded to the nearest whole

OpenStudy (anonymous):

\[\int \frac{1}{x^2(1-x)} dx\] Hmm I would like to try partial fraction \[\frac{1}{x^2(1+x)} = \frac{A}{1+x} + \frac{Bx +C}{x^2}\]

OpenStudy (anonymous):

partial fraction been a long time... i think its the denominator your working on right?

OpenStudy (anonymous):

\[1 = Ax^2 + Bx^2 + Bx + Cx + C \] Comparing co-efficient C = 1; (A+B) = 0; (B+C) = 0 Now you can integrate it! :-D

OpenStudy (anonymous):

Umm the Numerator \[\frac{1}{x^2(1+x)} = \frac{Ax^2 + Bx^2 + Bx + Cx + C}{x^2 (x+1)} = 1 = Ax^2 + Bx^2 + Bx + Cx + C\] Denominators cancel out.

OpenStudy (anonymous):

\[\frac{1}{x^2(1+x)} = \frac{Ax^2 + Bx^2 + Bx + Cx + C}{x^2 (x+1)} \implies 1 = Ax^2 + Bx^2 + Bx + Cx + C\] This version is much better

OpenStudy (anonymous):

your 1/x^2(1-x) becomes 1/x^2(1+x) typo?

OpenStudy (anonymous):

Oh Crap, yeah typo but the concept is same and I also messed up my calculations but the concept remains the same. Instead of \[\frac{1}{x^2(1+x)} = \frac{A}{1+x} + \frac{Bx +C}{x^2}\] you will have to do \[\frac{1}{x^2(1-x)} = \frac{A}{1-x} + \frac{Bx +C}{x^2}\]

OpenStudy (anonymous):

so its gonna be 1=\[x^2(A-B)+x(B-C)+C\]

OpenStudy (anonymous):

C= 1, (A-B) = 0, (B-C)=0

OpenStudy (anonymous):

yeah exactly!

OpenStudy (anonymous):

my answer is 50 =) is that right?